### v2

# A Reliable Framework for Human-in-the-Loop Anomaly Detection in Time Series

Abstract
Time series anomaly detection is a critical machine learning task for numerous applications, such as finance, healthcare, and industrial systems. However, even high-performed models may exhibit potential issues such as biases, leading to unreliable outcomes and misplaced confidence. While model explanation techniques, particularly visual explanations, offer valuable insights to detect such issues by elucidating model attributions of their decision, many limitations still exist—They are primarily instance-based and not scalable across dataset, and they provide one-directional information from the model to the human side, lacking a mechanism for users to address detected issues. To fulfill these gaps, we introduce HILAD, a novel framework designed to foster a dynamic and bidirectional collaboration between humans and AI for enhancing anomaly detection models in time series. Through our visual interface, HILAD empowers domain experts to detect, interpret, and correct unexpected model behaviors at scale. Our evaluation with two time series datasets and user studies demonstrates the effectiveness of HILAD in fostering a deeper human understanding, immediate corrective actions, and the reliability enhancement of models.
Index Terms:
Human-AI collaboration, time-series, anomaly detection, visual interface, XAI.
I Introduction
Time series anomaly detection is a crucial technique in machine learning (ML), with broad applications such as identifying irregular financial transactions [1], monitoring abnormal health indicators in patient records [2], and preempting equipment failures in manufacturing [3]. Despite its importance, the effectiveness of these models can be compromised by biases—where models make decisions based on incorrect input features, resulting in problematic conclusions. To address this, eXplainable AI (XAI) techniques, especially visual explanations, are employed to shed light on models’ decision-making processes. However, these methods often focus on specific instances without scaling or generalizing insights across different data groups. Moreover, they solely allow a one-way information flow from models to users, without enabling users to verify and act upon the derived insights. To address these challenges, we introduce HILAD, a human-in-the-loop framework to enhance the robustness and reliability of anomaly detection models. Through a carefully-designed interface, our framework empowers experts to effectively detect, interpret, and rectify model biases, thereby bolstering model performance and enhancing human trust. In the following, we delve deeper into different facets of our work.
I-A Time Series Anomaly Detection Models
ML techniques for time series anomaly detection can be broadly categorized into unsupervised and supervised approaches. When ground-truth anomaly labels are unknown, unsupervised learning techniques are introduced to approximate future data points [4, 5, 6], where the deviations between the predicted and true values serve as anomaly indicators. On the other hand, if labeled data are available, supervised models can be trained as classifiers to directly predict anomaly labels [7, 8, 9]. Such models can often achieve higher accuracy and be utilized in critical areas like fraud detection or medical diagnostics [10, 11]. However, supervised models are generally more susceptible to bias issues, causing significant risks [12]. Considering this, in our work, we focus on enhancing supervised models with a CNN architecture, aiming to detect, understand, and mitigate hidden biases in such models with a human-in-loop approach.
I-B XAI for Time Series Anomaly Detection
XAI methods are essential for enhancing transparency and building trust for time series anomaly detection models. Model agnostic approaches, such as LIME [13] and SHAP [14], attribute the model’s prediction of a specific instance to its features according to a surrogate model and game theory, respectively. Without the need to access models’ internal parameters, these methods are more generalizable for explaining various models, but fall short of specificity, tending to produce low-quality results, especially for deep models [15, 16, 17]. Contrastively, methods designed for specific model architectures can often provide more precise explanations with higher quality [18, 19, 20], despite compromising the flexibility to explain a broader range of models. For example, CAM and its variants [21, 22] are widely adopted in CNN explanations because of their capabilities of providing direct and intuitive explanations through heatmaps, highlighting which parts of the data mostly correspond to the model’s decision [23]. Specifically, CAM can assign a significance score to each time step of an input, indicating their temporal importance to a specific model prediction. Fig. 1 shows two time sequences with “anomaly” labels and a model makes “correct” predictions for both, where the actual periods that the anomaly happens are highlighted in blue and the model’s decision attributions provided by CAM are highlighted in green. In this example, it is clear that the blue and green regions overlap in (a) but not in (b), which means the model attributes anomaly prediction to normal time periods and generates seemingly correct results in (b). Example in (b) showcases the harmful model bias issue, and how CAM can detect such hidden issues.
Figure 1: Two sequences from the NASA MSL dataset with the ground truth label as “anomaly”. The true time frames corresponding to anomalies are highlighted in blue, and model’s attributes provided by CAM are highlighted in green. While the model makes “anomaly” prediction for both, regions highlighted by two colors overlap in (a) but not in (b), showing the model leverages the wrong features as the anomaly indicator in (b).
I-C Reliability Considerations of Models
Biased models can cause significant risks in real-world applications [24]. For instance, if a model identifies normal time series as anomalies, it underscores a fundamental misunderstanding of the underlying anomalous features, leading to potential overlooking of real anomalies [25]. While visual explanations such as CAM can reveal such discrepancies, their occurrence raises critical questions — Are these discrepancies a common feature across the entire data corpus, or are they isolated incidents? Moreover, once such biases are identified, what are the most effective methods for addressing and mitigating them? The discussion in Sec. I-B points out the limitations of current XAI techniques in fully tackling these challenges. This scenario necessitates further research and development of methods that not only highlight but also rectify these biases. Enhancing the reliability of anomaly detection models through improved understanding and correction of biases is essential to foster trust and ensure the efficacy of these systems in real-world applications.
I-D Potential of Human-AI Collaboration
Handling model reliability issues in time series anomaly detection is fundamentally a human-centric task. The complexity of determining what constitutes an anomaly necessitates a deep understanding that goes beyond algorithmic detection, requiring expert judgment and contextual awareness that machines alone are not equipped to provide. Researchers have actively developed visual analytics (VA) systems that incorporate human knowledge into the loop, designing visualization techniques to assist in the analysis of complex datasets and models in time series anomaly detection. However, current methods mainly involve single-instance inspections and do not provide effective summarization of predominant issues, thus limiting their ability to enhance models at scale. Additionally, these methods often lack sufficient interpretability support and require extensive manual verification by experts. To overcome these limitations, our work with HILAD aims to advance the state of the art by enhancing model interpretability, enabling the summarization and highlighting of critical data insights, thereby scaling up the integration of human knowledge into the machine learning process. HILAD is designed to streamline human-AI collaboration, making it more efficient at improving the accuracy and reliability of anomaly detection models.
I-E Our Contribution
HILAD is a framework with an interactive interface, designed to enhance the reliability of ML models in the context of classification-based anomaly detection models. Our framework leverages visual explanation techniques to detect model biases, integrated with attribution-aware clustering and visual information summary to support group-level model issue interpretation and validation. The interface of HILAD navigates domain experts through the model validation workflow, facilitating the identification and mitigation of potential model biases. Simultaneously, HILAD’s spuriousness estimation mechanism allows humans to contribute domain knowledge at scale, where they manually annotate a few problematic clusters and then verify the estimated spuriousness of others. Lastly, we present a method for enhancing the model by rectifying identified errors based on human feedback. This bidirectional interaction supported by our framework utilizes human expertise effectively with a feedback loop, allowing humans to enhance AI systems with minimal effort.
To demonstrate the effectiveness of HILAD, we invite domain experts to conduct model enhancement studies with benchmark datasets, and design comparative analyses to systematically compare our approach with state-of-the-art baselines through quantitative performance evaluation. Our evaluation encompasses both objective metrics and subjective assessments and aims to investigate the following hypotheses:
  1. [H1]
    Human-AI Collaborative Model Enhancement. The performance of the HILAD framework, when aligned with human’s domain expertise at the decision level, is significantly superior to conventional automated algorithms that operate on a black-box machine learning model.
  2. [H2]
    Human-Centric Model Interpretation. Participants’ subjective insights, in terms of intuitiveness, transparency, and reliability, will be significantly enhanced by direct interaction with models through HILAD framework. In contrast, approaches that lack this synergy may face challenges in offering such perceptions.
Our contributions are summarized as follows:
  1. HILAD, a framework that integrates human expertise with ML techniques to enhance the reliability of time-series anomaly detection models, allowing for systematically detecting, interpreting, and mitigating model biases.
  2. An interactive VA interface allowing for model validation at different scales, supporting effective issue annotation and propagation with minimal but essential human efforts.
  3. A solution with humans in the loop for generating actionable insights and incorporating such insights in model enhancement, which can inspire future research on human-assisted trustworthy AI.
II The Design of HILAD
Figure 2: The HILAD framework involves three main phases: “Behavior Summarization”, “Annotation and Propagation”, followed by “Model Enhancement”. With a time-series (TS) dataset and an anomaly detection model, we obtain model attributions with CAM. At the “Behavior Summarization” phase, we conduct attribution-aware clustering to obtain clusters with both similar TS data features and consistent model attributions, and we compute cluster summarization to aggregate each cluster’s pattern with a single visualization. At the “Annotation and Propagation” phase, users interact with our designed VA interface to identify and annotate model bias issues with the help of Spuriousness propagation. Lastly, the annotation and user-verified “Spuriousness score” are used at the “Model Enhancement” phase, which employs a model regularization method to mitigate detected model bias.
II-A HILAD Framework
HILAD accepts a time-series (TS) dataset with ground-truth anomaly labels and a trained ML model for anomaly classification as input; provides an interactive interface for domain experts to detect and understand model issues; and supports issue mitigation with an integrated model regularization technique. Fig. 2 illustrates the framework of HILAD. In our quest for a model behavior interpretation that is intuitive and has the potential to be summarized, we employ an XAI technique, CAM [21]. With individual time sequences and an ML classifier for anomaly detection, CAM can produce visual explanations highlighting model attributions of its decisions. In our “Behavior Summarization” phase, we compute the similarity metrics of the original TS data and the model attributions, respectively, to conduct attribution-aware clustering from the latent space, as detailed in Sec. II-C. To enhance the efficiency of identifying model’s biased attribution issues, we further summarize each cluster with an integrated visualization, enabling experts to verify the attribution correctness without inspecting individual sequences. Our following “Annotation and Propagation” phase involves a visual interface for user interactions, where users overview clusters’ summarized model attributions, validate the exhibition of model issues, and annotate verified issues with a few clicks. As long as at least one cluster with “spurious attention” has been annotated, our backend method propagates the user’s assessment across other clusters as the estimated “Spuriousness scores”, enabling a facilitated model validation. Detailed explanations of this process can be found in Sec. II-D. After identifying the incorrect model’s attention, our model regularization strategy at the “Model Enhancement” stage systematically addresses these inaccuracies, refining the model’s focus to improve its overall performance and reliability, as described in Sec. II-E.
Figure 3: HILAD applied to the anomaly detection of a univariate time series classifier trained on the MSL dataset [26]. A System Menu, enabling the selection of dataset, class, and visualization options (Scatter plot view or Detailed plot view). B Cluster Information window, showing the data clusters and cluster metrics. C Representation Space window, showing a visual overview and the relative positions of all data clusters. D Instance Information window, showing individual time series sequences with model attributions belonging to a selected data cluster.
II-B System Interface
The visual interface of HILAD comprises four main components. Fig. 3 depicts our interface when validating an anomaly classifier on the Mars Science Laboratory (MSL) dataset.
The System Menu (Fig. 3  A) provides options for configuration selections, including dataset, class, and visualization layout. Our primary objective is to analyze the model’s comprehension of anomalous patterns; therefore, we have configured the default class setting to represent anomaly-related instances, with ground-truth or prediction class as an anomaly. The Cluster Information (Fig. 3  B) presents an explainable behavior summarization for the current model in each cluster. Each cluster aggregates similar instances that garner analogous attention from the model. On the top of Cluster Information window, we introduce cluster metrics to facilitate navigation through the dataset. The available metrics, including Accuracy, Confidence, and Spuriousness — the latter derived via the Spuriousness Propagation method (Sec. II-E) contingent upon user annotation — are presented to enhance the rapidity of analysis. Once click/annotate one cluster in Cluster Information window, its corresponding position in 2D space will be highlighted in the Representation Space window (Fig. 3  C) and instances in this cluster will be detailed in the Instance Information window (Fig. 3  D). Two visualization layouts are available for Representation Space (Fig. 3  C): Scatter plot view or Detailed plot view, allowing for an overarching view of the similarity of clusters in 2D space or a more detailed inspection of each cluster. This function is also embedded in Representation Space to facilitate users to switch more flexibly. We also provide information on the correctness of the current model in terms of classification and attention.
Note that when interacting with the Representation Space window, the other two windows will be updated with coordinate information simultaneously to support user interpretation and issue verification. Due to the high similarity of adjacent clusters, when users annotate a cluster, they can quickly retrieve their neighbors in this window to speed up the annotation process. After a cluster is annotated in Cluster Information window, all instances of the cluster in Instance Information window have the same annotation by default. To further reduce the error introduced by the clustering algorithm, we allow users to adjust the annotations of individual data in the Instance Information window. The user must annotate data as correct or spurious at least once. Upon completing these annotations, the user can initiate model refinement by clicking the Retrain button. Subsequently, our backend algorithm will fine-tune the model in accordance with the user’s annotation, followed by an update to the model’s performance metrics shown in System Menu.
II-C Behavior Summarization
At the “Behavior Summarization” phase, we first apply CAM to obtain model attributions correlated to its decision for each individual sequence. Then, we measure the similarities between: (1) the time-series data, according to the Dynamic Time Warping (DTW) Metric, and (2) the model attribution masks, according to the cosine similarity, respectively. With this information, we perform attribution-aware clustering to obtain clusters with similar data features and model attributions. To allow for effective user investigation, we visually summarize each cluster to help users identify common patterns of each cluster at a simple glance. We describe the details in the following.
II-C1 Time-series Context Similarity Calculation
In the realm of time series analysis, assessing the similarity between sequences is a pivotal task. A prevalent technique to measure this similarity is the DTW algorithm, which strives to minimize the distance between two time-sequenced datasets by dynamically aligning and “warping” them in the time dimension. A smaller DTW distance indicates a higher similarity. Formally, for two multivariate time series $T_{Q}=\{{T_{q}}_{1},{T_{q}}_{2},\dots,{T_{q}}_{n}\}$ and $T_{P}=\{{T_{p}}_{1},{T_{p}}_{2},\dots,{T_{p}}_{m}\}$, ${T_{q}}_{i}\in\mathbb{R}^{d}$ and ${T_{p}}_{j}\in\mathbb{R}^{d}$ , the DTW distance $d_{DTW}(T_{Q},T_{P})$ can be computed using the following recursive formula:
$\begin{split}d_{DTW}(T_{Q},T_{P})=d({T_{q}}_{n},{T_{p}}_{m})\\ +\min\left\{\begin{array}[]{l}DTW({T_{q}}_{n-1},{T_{p}}_{m})\\ DTW({T_{q}}_{n},{T_{p}}_{m-1})\\ DTW({T_{q}}_{n-1},{T_{p}}_{m-1}),\end{array}\right.\end{split}$ (1)
where $d({T_{q}}_{n},{T_{p}}_{m})$ is the Euclidean distance between vectors ${T_{q}}_{n}$ and ${T_{p}}_{m}$. The initialization conditions are $D({T_{q}}_{n},{T_{p}}_{0})=D({T_{q}}_{0},{T_{p}}_{m})=\infty$ and $D({T_{q}}_{0},{T_{p}}_{0})=0$.
II-C2 Attribution Similarity Calculation
Given a fully convolutional network (FCN), $M$, trained for time series binary classification task (i.e., normal or abnormal), an input time series sequence, $T_{Q}$, and the predicted class of the $M$ for input $T_{Q}$, $c(T_{Q})$, the attention score of $T_{Q}$ is calculated as
$a(T_{Q})=CAM(M,T_{Q},c(T_{Q})),$ (2)
where $a(T_{Q})$ is a vector with the same dimension as $T_{Q}$, and each position in $a(T_{Q})$ corresponds to the degree of dependence of model $M$ on the corresponding dimension and time point when it makes the prediction $c(T_{Q})$. Thus $a(T_{Q})$ represents the model attribution. Please refer [21] to the specific calculation process of CAM. For input sequences $T_{Q}$ and $T_{P}$, the attribution similarity, $d_{cos}(a(T_{Q}),a(T_{P}))$, is computed using the cosine distance as:
$d_{cos}(a(T_{Q}),a(T_{P}))=1-\frac{a(T_{Q})\cdot a(T_{P})}{\|a(T_{Q})\|_{2}\times\|a(T_{P})\|_{2}},$ (3)
where $\|a(T_{Q})\|_{2}$ and $\|a(T_{P})\|_{2}$ denote the L2 norm of vectors $a(T_{Q})$ and $T_{a}(P)$, respectively. Consistent with $d_{DTW}$, a smaller cosine distance indicates a higher similarity.
II-C3 Attribution-Aware Clustering and Summarization
To aggregate data with similar context and also similar model attribution together, we construct the following distance $m_{QP}$ to measure the similarity between two instances $T_{Q}$ and $T_{P}$:
$m_{QP}=\alpha d_{DTW}(T_{Q},T_{P})+(1-\alpha)d_{cos}(a(T_{Q}),a(T_{P})),$ (4)
where $\alpha$ is the weight to balance the data context similarity $d_{DTW}(T_{Q},T_{P})$ and the model attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$. In this work, we set $\alpha=0.5$ in all case studies. Thus, for all of the $N$ time series instances, we can get the following distance matrix:
$M_{dist}=\begin{bmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N,1}&m_{N,2}&\cdots&m_{N,N}\end{bmatrix}.$ (5)
Finally, we use UMAP [27] to plot $M_{dist}$ and apply K-means [28] to generate $K$ clusters:
$C=\text{K-means}(UMAP(M_{dist}),K),$ (6)
where $C=\{c_{1},c_{2},\dots,c_{K}\}$ is the set of clusters, UMAP$(M_{dist})$ is the 2D representation of the data and $K$ is the desired number of clusters. As a follow-up step, we visually summarize the resulting clusters as an integrated visualization for each cluster, which presents the common patterns of time sequences and model attributions in the current cluster, empowering users to grasp model behavior patterns more efficiently.
II-D Annotation and Propagation
Building upon the aforementioned method, at the “Annotation and Propagation” phase, we To effectively evaluate such spurious attention, we introduce a metric termed “Spuriousness score”, $S[c_{i}]$, ranging from 0 to 1, signifying the probability range from the least to the most likelihood of spurious attention occurrence. Furthermore, we deploy the Label Propagation [29], which automatically estimates probabilities of the unannotated clusters’ spuriousness based on users’ annotations and neighboring clusters $N(c_{i})$ in the representation space.
The representation space is generated by UMAP as described in Sec. II-C, where we use the adjacency matrix, $A(i,j)$, to denote the distance between clusters’ feature representations of $c_{i}$ and $c_{j}$. Notably, users only need to provide an annotation for one cluster (Algo. 1, Line 6), designating it as either “correct attention” ($S[c_{i}]=0$) or “spurious attention” ($S[c_{i}]=1$). Our algorithm (Algo. 1, Line 8-15) then propagates this information to compute Spuriousness scores for other clusters. Subsequently, these scores are showcased in the Cluster Information window (Fig. 3  B).
The “Spuriousness score” yields two noteworthy advantages. Primarily, they facilitate a more streamlined exploration process of clusters, with potentially spurious correlations being conspicuously highlighted. This is a pivotal step in aiding users to identify and assess problematic attention. Secondly, following users’ validation, these scores are leveraged to determine the problematic clusters crucial for issue mitigation, as we elucidate in the following section.
Algorithm 1 Label Propagation for Spuriousness Score
1:
2:$C$ : Set of clusters.
3:$A$ : Adjacency matrix of clusters indicating similarity.
4:
5:$\mathbf{S}$ : Vector of spuriousness scores for each cluster.
6:Initialize $\mathbf{S}$ with -1 for all clusters.
7:for each cluster $c_{i}$ in $C$ with a user label do
8:     $S[c_{i}]$ = label value of $c_{i}$
9:while unlabeled clusters exist in $C$ do
10:     for each unlabeled cluster $c_{i}$ in $C$ do
11:         $\text{WeightedSum}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})\times S[c_{j}]$
12:         $\text{TotalWeight}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})$
13:         $S[c_{i}]=\frac{\text{WeightedSum}(c_{i})}{\text{TotalWeight}(c_{i})}$
14:return $\mathbf{S}$
II-E Model Enhancement
An effective strategy for addressing detected spurious correlations in machine learning (ML) models involves model retraining or finetuning, which reduces the model’s resilience over potential bias while keeping the architecture intact. Retraining the model on a noise-corrupted dataset has demonstrated its efficacy in diminishing the model’s reliance on spurious features [30]. To counteract the influence of spurious features, we harness the masking strategy to shift and correct the model’s attention. The idea is conceptually simple: first, mask the parts of the data that receive spurious attention, then encourage the model to infer how to make decisions with the rest of the data. This idea has been used as a self-supervised pre-training method and spuriousness feature elimination method in Computer Vision (CV) and Natural Language Processing (NLP) fields. In our work, we adjust the Core Risk Minimization (CoRM) methodology [30] in the CV domain into time series data to introduce controlled masks into the annotated spurious attention time points by injecting random Gaussian noise into the data. This is also a fundamental data augmentation strategy called jittering for time series data that can help the model become more robust to variations and be used in masked autoencoder for time series forecasting task [31].
The integration of CoRM into HILAD entails several steps. First, after users annotate the data, we extract a set of instances, $T_{S}$, that display heightened probabilities of spuriousness, indicating undesirable attention. We also extract a set of instances, $T_{C}$, that display heightened probabilities of correctness, indicating correct attention. For the instances in $T_{S}$, pixel attributions (specifically, CAM masks) accentuate the spurious regions. We employ these masks to introduce random Gaussian noise to the identified spurious regions. Mathematically, for an individual time sequence $T_{i}\in T_{S}$, this operation can be succinctly represented as
$T^{\prime}_{i}=T_{i}+\mathbf{m}\odot z,$ (7)
wherein $T_{i}$ is the input time series data, $\mathbf{m}$ denotes the CAM mask, $z$ represents the generated Gaussian noise matrix, and $\odot$ denotes the Hadamard product. Each variable is of the same dimensions as the input time series data. Similarly, for $T_{j}\in T_{C}$, we add random Gaussian noise to the non-highlighted areas to enhance the model’s attention to abnormal patterns:
$T^{\prime}_{j}=T_{j}+(1-\mathbf{m})\odot z.$ (8)
After substituting the original data with these spuriousness-masked and correctness-enhanced data, we proceed to retrain the model. Subsequently, an evaluation is conducted to determine the extent to which spurious correlations have been alleviated. In Sec. III-D, we elaborate on the evaluation metrics utilized to quantify the effectiveness of our method in mitigating spurious correlations.
III Experimental Design
In this section, we provide an experimental design to benchmark and evaluate the capabilities of HILAD. The primary objective of this experiment is to investigate how HILAD empowers humans to detect, understand, and correct potential issues in an anomaly classifier for time series data.
III-A Task
In the experiment, participants were asked to interact with the interface of HILAD to evaluate a classification model trained on a specific dataset, containing both normal and abnormal time sequences. Specifically, we encouraged them to use as many visual components as possible, follow the system guidance to verify the highlighted issues, interpret them with the provided information, and annotate the confirmed errors in a restricted time. In each experiment, participants were only allowed to perform a single model retraining to better control experiment time and quantify the evaluation results.
With the awareness of the HILAD workflow, participants explored the general model behavior concerning different clusters of data through the Cluster Information window, where they also investigated further details from the Instance Information window that updated simultaneously with their interactions. They were asked to mark and annotate detection issues, such as a model using normal time points to determine an abnormal time series, which could be achieved by simply clicking the “Spurious” button in the Cluster Information window. At this point, all instances corresponding to this cluster were marked as spurious. Participants could further adjust the spuriousness labels of single instances in the Instance Information window. Similarly procedure could be conducted to label and adjust a cluster related to correct attention. As long as one cluster is annotated as spurious or correct, our algorithm will approximate the attention correctness of other clusters according to their similarities in the representation space, resulting in a spuriousness approximation score for every cluster. Participants can click the “spuriousness” button provided in our interface to rank clusters according to this score, allowing for adjustments and confirmation of the approximated model issues. When the participants confirmed the first round of issue inspection could be concluded, they would click the retrain button, and our system performed the model retraining based on the participants’ feedback.
III-B Procedure
In the following, we introduce the experiment’s procedure. The overall practical accomplishment of the experiment took between 30 and 45 minutes.
III-B1 Introduction and Preparations
Before the experiment begins, participants receive detailed instructions explaining the experiment’s purpose and overall structure before it begins. Next, they are introduced to the key concepts involved in time-series classification tasks, including metrics such as confidence and accuracy, possible abnormal patterns, and spurious model attention. Subsequently, an in-depth orientation to HILAD is provided, highlighting its various components and functionalities. This orientation covers the System Menu, Representation Space, Cluster Information, and Instance Information, elucidating their respective roles in detecting anomalies. The primary goal communicated to the participants is to employ HILAD for monitoring effectively and analyzing the model’s behavior, focusing on its attention mechanisms.
III-B2 Familiarization Procedure
A simplified dataset was introduced to acquaint participants with the functions of the HILAD interface. This served as a practical tutorial on the tools available in each window. During this phase, participants gained insights into the relationships between different clusters and developed skills in identifying and annotating abnormal patterns. It is important to note that this segment was designed solely for hands-on learning, and its results were not included in the primary evaluation. Once participants demonstrated proficiency in accurately identifying and annotating abnormal patterns, they conducted the experimental run.
III-B3 Experimental Run
In the experimental run, participants were introduced to the HILAD with two space domain-related datasets that mimic its real-world use for anomaly detection. Their task was to critically analyze the behavior of the classification model across the datasets. The main variable in this phase was the dataset itself, and the sequence of dataset presentation varied systematically among participants. This approach aimed to mitigate any potential bias from learning curve effects. We executed three trials for each dataset to further effectively compare the distinctions between the automated algorithm and the human-AI collaboration method. These trials varied in the scope of information provided to the participants: the first trial provided access to only the System Menu window, the second trial included System Menu, Cluster Information, and Representation Space windows, and the third trial offered a comprehensive view with the complete all of the four windows. The first trial in our study exemplifies the application of the automated algorithm, while the latter two trials represent the human-AI collaboration methods. To ensure the integrity of our results and eliminate any potential biases associated with the sequence of exposure, the order of these trials was randomized for each participant. This approach was essential to assess the impact of varying levels of information availability on the performance and decision-making processes in the context of human-AI interaction.
III-B4 Postprocessing
Once all experimental runs are concluded, participants can reflect on the experiment and share their insights and feedback regarding HILAD. This qualitative data is gathered through questionnaires designed to assess participants’ subjective perceptions of human-AI cooperation, specifically focusing on intuition, transparency, and reliability. This data provides valuable insights into the user experience and the system’s effectiveness in facilitating anomaly detection, evaluation, and correction.
III-C Participants
This experiment was approved by the Institutional Review Board at the University of California, Davis. We involve 15 participants in total, of which 13 were male and 2 were female. All participants are graduate students between the ages of 23 to 30 years old. All participants passed the test to ensure they could accurately identify and annotate the abnormal pattern. All completed the entire experiment on two provided datasets.
III-D Performance Evaluation Metrics
We measure the effectiveness of our proposed framework in terms of classification and attention performance. From the perspective of classification performance, we use accuracy ($Acc$), precision ($P$), recall ($R$), and $F_{1}$ score to evaluate the performance of our method comprehensively. The details of the metrics are listed as follows:
$Acc=\frac{TP+TN}{TP+TN+FP+FN},P=\frac{TP}{TP+FP},$
$R=\frac{TP}{TP+FN},F_{1}=\frac{2\cdot P\cdot R}{P+R},$
where TP, TN, FP, and FN are four classification results, denoting true positive, true negative, false positive, and false negative, respectively. $Acc$ is a measure of the overall correctness of a classification model. $P$ represents the ability of the method to distinguish anomalies, and a higher $P$ accurately indicates fewer false anomalies. $R$ is to determine the ability of the method to detect all the anomalies. Since The $P$ and $R$ metrics often appear contradictory, the $F_{1}$ score is a metric for comprehensively considering these two indicators and providing a balanced overall performance of the method. When the data set has imbalanced classes, F1 Score might be a better metric than accuracy.
In the perspective of attention accuracy performance, we adjust the relevance accuracy ($RA$) proposed in [23] to quantify how accurate a model is at locating ground-truth time points at which the abnormal behavior occurs. The core idea of relevance accuracy is to compare the $k$ most attention time points for a prediction with the ground-truth abnormal time points. For a given time sequence $T_{i}$, let $G(T_{i})$ denote the set of continuous or discontinuous ground-truth abnormal time points, where $|G(T_{i})|$ is the number of ground-truth time points, and $R_{i}(k)$ denote the $k$ time steps that get the most attention for the prediction of the model. To apply this metric to data with varying numbers of ground-truth time points, we set $k=(1+M\%)\times|G(T_{i})|$ and denote $R_{i}(k)$ as $R_{i}(M\%)$ in our work. In this analysis, the sequence of the ground-truth time points is disregarded, given that each ground-truth time point is treated with equal significance. Optimal relevance accuracy is attained when all elements of $G(T_{i})$ are encompassed within $R_{i}(M\%)$, ideally with the minimal possible value of $M$. For a time sequence $T_{i}$, $RA(T_{i})$ can be calculated as:
$RA(T_{i})=\frac{|R_{i}(M\%)|\cap G(T_{i})}{|G(T_{i})|},$ (9)
where $\cap$ is the intersection of two sets and ||$\cdot$|| ⋅ | is the cardinality of a set. The time points the model is attention to are ranked based on CAM value before they are compared to $G(T_{i})$. Therefore, simply highlighting all time points will not result in a high relevance accuracy score. The model needs to highlight some time points as being more important than others.
To assess the participants’ subjective perception of human–machine cooperation, a questionnaire with a five-point Likert Scale [24] with the following items was used.
Figure 4: Tukey HSD post-hoc test results for MSL dataset. The left plot illustrates classification performance, while the right one depicts attention accuracy. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
III-E Datasets
Mars Science Laboratory (MSL) dataset is a dataset that corresponds to the sensor and actuator data for the Mars rover [26]. At the same time, the data set marks the time point when the abnormal behavior occurs for us to calculate relevance accuracy. However, this dataset is known to have many trivial sequences [32]; hence, to make a best-faith effort to filter out data with mislabeled ground truth time points, we build the new dataset based on the two nontrivial ones (A2 and A4), where we randomly divide the original time series into 1000 equal-length sequences, all of 800 lengths, where the ratio of labeled normal to abnormal data is 3 to 2.
The Simulation Testbed for Exploration Vehicle ECLSS (STEVE) [33] at the University of Colorado Boulder is a simplified single-bed $CO_{2}$ removal system of the Carbon Dioxide Removal Assembly (CDRA) onboard the International Space Station (ISS). A corresponding Simulink model is provided to simulate multiple failure modes similar to the STEVE testbed. In our study, we conducted a simulation consisting of four cycles, each comprising 80 minutes of $CO_{2}$ adsorption followed by 80 minutes of desorption. During the third cycle, a simulated leak failure is introduced at a random time point in Valve 1, resulting in observable anomalies in three sensors: Bed Temperature, $CO_{2}$ Concentration, and Flow Rate. These sensor sequences are treated collectively as a set of sequences. We generated 1000 sets of sequences with a ratio of 4 to 1 between labeled normal and abnormal data. Additionally, we marked the time points at which abnormal behaviors occurred in the three sensors to calculate relevance accuracy.
III-F Baseline Methods
Baseline model. In this paper, we use the fully convolutional network (FCN) proposed in [34], which has demonstrated good performance on time series classification benchmarks. In addition to the baseline model, we also compare our spurious attention elimination method with the following data augmentation strategies for the same number, $Q$, of instances:
Random Aug. Randomly select $Q$ instances from the training set for model fine-tuning.
Random Mask. Select the $Q$ instances proposed by our system, but add masks at random time points for each instance. To further verify the effectiveness of our method, we compared the following strategy as the ablation study of our method.
Cluster Mask. The $Q$ instances proposed by our system only concede the participant’s operation in the Cluster Information window, which means the participants in the Instance Information window make no further adjustments. This method uses the same masking strategy as ours.
III-G Statistical Analysis
For Hypothesis 1, since all five methods involved in the comparison have been tested an equal number of times, we directly assess their performance using the metrics introduced in Section III-D. Therefore, our hypothesis tests reduces to:
$H_{0}:{\mu_{Acc}}_{1}=\dots={\mu_{Acc}}_{5},{\mu_{P}}_{1}=\dots={\mu_{P}}_{5},$
${\mu_{R}}_{1}=\dots={\mu_{R}}_{5},{\mu_{F_{1}}}_{1}=\dots={\mu_{F_{1}}}_{5},$
${\mu_{RA}}_{1}=\dots={\mu_{RA}}_{5};$
$H_{1}:{\mu_{Acc}}_{1}\neq\dots\neq{\mu_{Acc}}_{5},{\mu_{P}}_{1}\neq\dots\neq{\mu_{P}}_{5},$
${\mu_{R}}_{1}\neq\dots\neq{\mu_{R}}_{5},{\mu_{F_{1}}}_{1}\neq\dots\neq{\mu_{F_{1}}}_{5},$
${\mu_{RA}}_{1}\neq\dots\neq{\mu_{RA}}_{5}.$
To evaluate participants’ subjective perceptions of human-AI cooperation, we administered a questionnaire featuring a five-point Likert scale designed to assess key aspects such as intuition, transparency, and reliability. In this subjective evaluation, we compare three methods: (1) Random Augmentation, driven by a purely automatic algorithm. When employing this method, only the System Menu is accessible within our interactive interface, and the user interaction is limited to operating the retrain button; (2) Human-AI Cluster Mask, which, upon execution, presents windows A, B, and C on our interface. Users can assess the correctness of attention by clusters but lack the capability to inspect and adjust individual data points within each cluster; (3) Human-AI HILAD, the comprehensive version, where users are allowed to refine their assessments in window D. By contrasting (1) with (2) and (3), we aim to elucidate the impact of increasing user engagement on subjective evaluation. Conversely, by contrasting (2) with (3), we seek to discern the added benefits by enabling detailed individual data adjustments within the clusters.
IV Results
Figure 5: Tukey HSD post-hoc test results for STEVE dataset. The left plot illustrates classification performance, while the right one depicts attention accuracy. Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
Figure 6: Participants perception of the system, according to nine Likert-Type Questions using a 5-point scale.
In order to effectively measure the objective performance difference between purely automatic algorithm-driven and human-machine cooperative anomaly detection methods, we conducted an analysis using Analysis of Variance (ANOVA) tests [35]. Objective performance is divided into two aspects, namely classification correctness and attention correctness. For the former, we utilize the ANOVA test to compare the means of accuracy, precision, recall, and F1 score across different models to statistically ascertain any significant performance variations. For the latter, we employed the ANOVA test to determine whether there are statistically significant differences in relevance accuracy at varying levels of $M$, specifically at $M=20,25,30$, and $35$. Upon identifying significant differences with the ANOVA test, we proceeded with the Tukey HSD post-hoc test to further distinguish which specific group means, among the compared groups, are significantly different from each other, ensuring a thorough evaluation of pairwise contrasts. Owing to space limitations, this article presents only the results of the Tukey HSD post-hoc test for two datasets, as illustrated in Fig. 4 and Fig. 5, respectively.
The result of the five-point Likert scale for objective evaluation is shown in Fig. 6. For a more comprehensive analysis, we also furnish the post-test results for the three methods, assessing their comparative effectiveness using a t-test. The stars intended to flag levels of significance as shown in Fig. 6 are based on the adjusted p-values.
V Discussion
V-A Hypothesis 1
Firstly, regarding classification correctness, Fig. 4 demonstrates that on univariate data, all data augmentation methods significantly enhance the model’s classification performance compared to the baseline model. Among these methods, within the context of human-AI cooperation, the Cluster Mask approach does not exhibit a significant difference from the Random Mask in the automatic method across other metrics, with the exception of the F1 score. However, our method demonstrated significant enhancements in all four metrics of classification performance when compared to all other methods. Compared to approach Cluster Mask, the improvement in our approach is facilitated by the   D window’s functionality, which enables humans to further refine the automatic clustering algorithm. This aspect also contributes to the observed large variance in our method across the four metrics, reflecting the variability in human judgment and adjustments. This variability is partly due to the high number of instances in each cluster. Despite this, in terms of overall performance, the adjustments made by humans do yield beneficial outcomes.
As depicted in Fig. 5, similar to the univariate data, all data augmentation methods, with the exception of the accuracy metric for the Random Mask approach, show results that are significantly different from the baseline model. Furthermore, within the realm of human-AI cooperation, the Cluster Mask approach exhibits a significant difference over other automatic algorithms solely in terms of accuracy. However, our method demonstrates significant differences from both other automatic data augmentation methods and the Cluster Mask method in terms of accuracy, precision, and F1 score. In terms of recall, no significant differences are observed among all data augmentation methods. This phenomenon can be attributed to the imbalanced nature of the dataset, where the abundance of normal instances prompts the model to occasionally misclassify abnormal instances as normal. In such imbalanced datasets, neither recall nor precision in isolation serves as a reliable measure of method performance. Instead, the F1 score emerges as a more accurate metric due to its ability to strike a balance between precision and recall.
Shifting our focus to attention accuracy, both the cluster mask and our method exhibit significant deviations from other automatic algorithms when applied to univariate data. However, as the data dimensionality increases, achieving accurate attention becomes increasingly challenging, particularly at $M=20$, where no significant disparities exist between the cluster mask method and other automatic algorithms. Nevertheless, consistent disparities persist between our methods and the automatic algorithm. A more in-depth analysis reveals that, despite the notable improvements in classification performance when compared to the baseline model, the automatic algorithm fails to significantly enhance attention accuracy, especially concerning multivariate data. This observation reinforces the notion that our method not only objectively enhances classification performance but also elevates attention accuracy, thereby bolstering overall system reliability.
In summary, our method exhibits substantial deviations in both classification correctness and attention correctness when compared to the automatic algorithm across the two datasets. This substantiates our conjecture and underscores the empirical evidence supporting our approach.
V-B Hypothesis 2
In the subjective evaluation, both the cluster mask method and our method are significantly different from the automatic algorithm in intuition, transparency, and reliability. Because the automatic algorithm can only provide the result of classification, can not provide the criterion of classification can not interact. The cluster mask method is not significantly different from intuition and transparency, as both provide reasons for the model to make a prediction. But on reliability, we are significantly different from the cluster mask method because of the error of the automatic clustering algorithm itself, while in our method, the D window can provide more detailed options to further correct the automatic algorithm, which helps improve performance objectively. At the same time, humans are subjectively more likely to believe that our methods are more reliable.
Therefore, we verified that subjectively human-AI method has significant improvement in intuition, transparency, and reliability compared with automatic algorithms, thus verifying Hypothesis 2. Further, on the objective and subjective level, the necessity of introducing D window is verified by comparing with Cluster Mask method.
VI Conclusion
In conclusion, we introduce a human-AI collaboration framework, HILAD, to enhance the performance and reliability of time series anomaly detection models. With our interactive visual interface, human expertise can be effectively leveraged to detect and eliminate the hidden model biases in scale. Through objective and subjective evaluations with two benchmark datasets, we demonstrate the exceptional capabilities of HILAD in enhancing the model’s classification accuracy and attention correctness. With HILAD, we underscore the potential of human-AI collaboration in supporting greater transparency, reliability, and trustworthiness of machine learning models. We hope the design of how human expertise can be effectively leveraged in our work can inspire the development of next-generation trustworthy AI systems.

### v4

@@
 Time series anomaly detection is a critical machine learning task for numerous applications, such as finance, healthcare, and industrial systems.
-However, even high-performed models may exhibit potential issues such as biases, leading to unreliable outcomes and misplaced confidence.
-While model explanation techniques, particularly visual explanations, offer valuable insights to detect such issues by elucidating model attributions of their decision, many limitations still exist—They are primarily instance-based and not scalable across dataset, and they provide one-directional information from the model to the human side, lacking a mechanism for users to address detected issues.
+However, even high-performing models may exhibit potential issues such as biases, leading to unreliable outcomes and misplaced confidence.
+While model explanation techniques, particularly visual explanations, offer valuable insights by elucidating model attributions of their decision, many limitations still exist—They are primarily instance-based and not scalable across the dataset, and they provide one-directional information from the model to the human side, lacking a mechanism for users to address detected issues.
 To fulfill these gaps, we introduce HILAD, a novel framework designed to foster a dynamic and bidirectional collaboration between humans and AI for enhancing anomaly detection models in time series.
 Through our visual interface, HILAD empowers domain experts to detect, interpret, and correct unexpected model behaviors at scale.
-Our evaluation with two time series datasets and user studies demonstrates the effectiveness of HILAD in fostering a deeper human understanding, immediate corrective actions, and the reliability enhancement of models.
-Index Terms:
-Human-AI collaboration, time-series, anomaly detection, visual interface, XAI.
-I Introduction
-Time series anomaly detection is a crucial technique in machine learning (ML), with broad applications such as identifying irregular financial transactions [1], monitoring abnormal health indicators in patient records [2], and preempting equipment failures in manufacturing [3].
+Our evaluation through user studies with two models and three time series datasets demonstrates the effectiveness of HILAD, which fosters a deeper model understanding, immediate corrective actions, and model reliability enhancement.
+1 Introduction
+Time series anomaly detection is a crucial technique in machine learning (ML), with broad applications such as identifying irregular financial transactions Ahmed et al.
+(2016), monitoring abnormal health indicators in patient records Šabić et al.
+(2021), and preempting equipment failures in manufacturing Li et al.
+(2019).
 Despite its importance, the effectiveness of these models can be compromised by biases—where models make decisions based on incorrect input features, resulting in problematic conclusions.
-To address this, eXplainable AI (XAI) techniques, especially visual explanations, are employed to shed light on models’ decision-making processes.
-However, these methods often focus on specific instances without scaling or generalizing insights across different data groups.
-Moreover, they solely allow a one-way information flow from models to users, without enabling users to verify and act upon the derived insights.
-To address these challenges, we introduce HILAD, a human-in-the-loop framework to enhance the robustness and reliability of anomaly detection models.
-Through a carefully-designed interface, our framework empowers experts to effectively detect, interpret, and rectify model biases, thereby bolstering model performance and enhancing human trust.
-In the following, we delve deeper into different facets of our work.
-I-A Time Series Anomaly Detection Models
-ML techniques for time series anomaly detection can be broadly categorized into unsupervised and supervised approaches.
-When ground-truth anomaly labels are unknown, unsupervised learning techniques are introduced to approximate future data points [4, 5, 6], where the deviations between the predicted and true values serve as anomaly indicators.
-On the other hand, if labeled data are available, supervised models can be trained as classifiers to directly predict anomaly labels [7, 8, 9].
-Such models can often achieve higher accuracy and be utilized in critical areas like fraud detection or medical diagnostics [10, 11].
-However, supervised models are generally more susceptible to bias issues, causing significant risks [12].
-Considering this, in our work, we focus on enhancing supervised models with a CNN architecture, aiming to detect, understand, and mitigate hidden biases in such models with a human-in-loop approach.
-I-B XAI for Time Series Anomaly Detection
-XAI methods are essential for enhancing transparency and building trust for time series anomaly detection models.
-Model agnostic approaches, such as LIME [13] and SHAP [14], attribute the model’s prediction of a specific instance to its features according to a surrogate model and game theory, respectively.
-Without the need to access models’ internal parameters, these methods are more generalizable for explaining various models, but fall short of specificity, tending to produce low-quality results, especially for deep models [15, 16, 17].
-Contrastively, methods designed for specific model architectures can often provide more precise explanations with higher quality [18, 19, 20], despite compromising the flexibility to explain a broader range of models.
-For example, CAM and its variants [21, 22] are widely adopted in CNN explanations because of their capabilities of providing direct and intuitive explanations through heatmaps, highlighting which parts of the data mostly correspond to the model’s decision [23].
-Specifically, CAM can assign a significance score to each time step of an input, indicating their temporal importance to a specific model prediction.
-Fig.
-1 shows two time sequences with “anomaly” labels and a model makes “correct” predictions for both, where the actual periods that the anomaly happens are highlighted in blue and the model’s decision attributions provided by CAM are highlighted in green.
-In this example, it is clear that the blue and green regions overlap in (a) but not in (b), which means the model attributes anomaly prediction to normal time periods and generates seemingly correct results in (b).
-Example in (b) showcases the harmful model bias issue, and how CAM can detect such hidden issues.
+To tackle this issue, explainable AI (XAI) techniques Dwivedi et al.
+(2023); Xuan et al.
+(2024a) have been introduced to expose the decision logic of anomaly detectors.
+Among them, visual explanation methods like Class Activation Mapping (CAM) Zhou et al.
+(2016); Boniol et al.
+(2023) are particularly effective in attributing importance scores to each time step in a sequence.
+CAM can assign a significance score to each input time step, indicating its contribution to a model prediction.
+For example, Fig.
+1 illustrates two time sequences with both labeled as “anomaly”, and the model makes correct predictions.
+However, CAM reveals a critical distinction: in case (a), the model’s attention (green) aligns well with the actual anomalous segment (blue), while for (b), the attribution of the prediction (green) does not overlap with the ground-truth (blue).
+Although both predicted labels seem correct, only the first reflects proper reasoning.
+This discrepancy demonstrates a hidden model bias, where decisions appear valid but are based on faulty logic, and highlights the need for an explainable system that not only exposes the bias but also helps mitigate such failures.
+However, most existing XAI methods are limited in two key aspects: they analyze individual instances without generalizing to systematic model behaviors, and they enable only one-way interpretation—from model to user—without allowing users to intervene or improve the model Barkouki et al.
+(2023).
+In complex real-world environments, such limitations can lead to brittle deployments, where spurious model logic remains undetected and uncorrected at scale.
+To ensure robust and trustworthy anomaly detection, we argue that explanation alone is insufficient.
+Instead, effective systems must support human-in-the-loop workflows where users can not only validate model decisions but also provide corrective feedback that influences model behavior.
+This perspective is supported by recent progress in human-AI collaboration, which shows how visual analytics interfaces can empower users to audit, refine, and debug AI models Amershi et al.
+(2019); Xuan et al.
+(2025a, b).
 Figure 1: Two sequences from the NASA MSL dataset with the ground truth label as “anomaly”.
-The true time frames corresponding to anomalies are highlighted in blue, and model’s attributes provided by CAM are highlighted in green.
-While the model makes “anomaly” prediction for both, regions highlighted by two colors overlap in (a) but not in (b), showing the model leverages the wrong features as the anomaly indicator in (b).
-I-C Reliability Considerations of Models
-Biased models can cause significant risks in real-world applications [24].
-For instance, if a model identifies normal time series as anomalies, it underscores a fundamental misunderstanding of the underlying anomalous features, leading to potential overlooking of real anomalies [25].
-While visual explanations such as CAM can reveal such discrepancies, their occurrence raises critical questions — Are these discrepancies a common feature across the entire data corpus, or are they isolated incidents?
-Moreover, once such biases are identified, what are the most effective methods for addressing and mitigating them?
-The discussion in Sec.
-I-B points out the limitations of current XAI techniques in fully tackling these challenges.
-This scenario necessitates further research and development of methods that not only highlight but also rectify these biases.
-Enhancing the reliability of anomaly detection models through improved understanding and correction of biases is essential to foster trust and ensure the efficacy of these systems in real-world applications.
-I-D Potential of Human-AI Collaboration
-Handling model reliability issues in time series anomaly detection is fundamentally a human-centric task.
-The complexity of determining what constitutes an anomaly necessitates a deep understanding that goes beyond algorithmic detection, requiring expert judgment and contextual awareness that machines alone are not equipped to provide.
-Researchers have actively developed visual analytics (VA) systems that incorporate human knowledge into the loop, designing visualization techniques to assist in the analysis of complex datasets and models in time series anomaly detection.
-However, current methods mainly involve single-instance inspections and do not provide effective summarization of predominant issues, thus limiting their ability to enhance models at scale.
-Additionally, these methods often lack sufficient interpretability support and require extensive manual verification by experts.
-To overcome these limitations, our work with HILAD aims to advance the state of the art by enhancing model interpretability, enabling the summarization and highlighting of critical data insights, thereby scaling up the integration of human knowledge into the machine learning process.
-HILAD is designed to streamline human-AI collaboration, making it more efficient at improving the accuracy and reliability of anomaly detection models.
-I-E Our Contribution
-HILAD is a framework with an interactive interface, designed to enhance the reliability of ML models in the context of classification-based anomaly detection models.
-Our framework leverages visual explanation techniques to detect model biases, integrated with attribution-aware clustering and visual information summary to support group-level model issue interpretation and validation.
-The interface of HILAD navigates domain experts through the model validation workflow, facilitating the identification and mitigation of potential model biases.
-Simultaneously, HILAD’s spuriousness estimation mechanism allows humans to contribute domain knowledge at scale, where they manually annotate a few problematic clusters and then verify the estimated spuriousness of others.
+The true time frames corresponding to anomalies are highlighted in blue, and model’s CAM attributes are highlighted in green.
+While the model makes “anomaly” prediction for both, regions highlighted by two colors overlap in (a) but not in (b), showing the model leverages wrong features as the anomaly indicator in (b).
+In this paper, we introduce HILAD, a human-in-the-loop framework with an interactive interface, designed to enhance the reliability of ML models in the context of time series anomaly classification.
+Our framework leverages visual explanation techniques to detect model biases, designing attribution-aware clustering and summarization to foster model issue interpretation and validation.
+The interface of HILAD navigates domain experts through the model validation workflow, facilitating the identification and mitigation of model biases.
+Our spuriousness propagation mechanism allows humans to contribute domain knowledge at scale, where they only need to manually annotate a few clusters, and the estimated spuriousness scores of others are provided automatically for their verification.
 Lastly, we present a method for enhancing the model by rectifying identified errors based on human feedback.
-This bidirectional interaction supported by our framework utilizes human expertise effectively with a feedback loop, allowing humans to enhance AI systems with minimal effort.
-To demonstrate the effectiveness of HILAD, we invite domain experts to conduct model enhancement studies with benchmark datasets, and design comparative analyses to systematically compare our approach with state-of-the-art baselines through quantitative performance evaluation.
+Our bidirectional interaction utilizes human expertise effectively with a feedback loop, allowing humans to enhance AI systems with minimal effort.
+To demonstrate the effectiveness of HILAD, we invite domain experts to conduct model enhancement studies with benchmark datasets, and design comparative analyses to quantitatively compare our approach with representative baselines.
 Our evaluation encompasses both objective metrics and subjective assessments and aims to investigate the following hypotheses:
@@
   1.
-HILAD, a framework that integrates human expertise with ML techniques to enhance the reliability of time-series anomaly detection models, allowing for systematically detecting, interpreting, and mitigating model biases.
+HILAD, a framework that integrates human expertise with ML techniques to enhance the reliability of time series anomaly detection models, allowing for systematically detecting, interpreting, and mitigating model biases.
   2. An interactive VA interface allowing for model validation at different scales, supporting effective issue annotation and propagation with minimal but essential human efforts.
-  3. A solution with humans in the loop for generating actionable insights and incorporating such insights in model enhancement, which can inspire future research on human-assisted trustworthy AI.
-II The Design of HILAD
-Figure 2: The HILAD framework involves three main phases: “Behavior Summarization”, “Annotation and Propagation”, followed by “Model Enhancement”.
-With a time-series (TS) dataset and an anomaly detection model, we obtain model attributions with CAM.
-At the “Behavior Summarization” phase, we conduct attribution-aware clustering to obtain clusters with both similar TS data features and consistent model attributions, and we compute cluster summarization to aggregate each cluster’s pattern with a single visualization.
-At the “Annotation and Propagation” phase, users interact with our designed VA interface to identify and annotate model bias issues with the help of Spuriousness propagation.
-Lastly, the annotation and user-verified “Spuriousness score” are used at the “Model Enhancement” phase, which employs a model regularization method to mitigate detected model bias.
-II-A HILAD Framework
-HILAD accepts a time-series (TS) dataset with ground-truth anomaly labels and a trained ML model for anomaly classification as input; provides an interactive interface for domain experts to detect and understand model issues; and supports issue mitigation with an integrated model regularization technique.
-Fig.
-2 illustrates the framework of HILAD.
-In our quest for a model behavior interpretation that is intuitive and has the potential to be summarized, we employ an XAI technique, CAM [21].
-With individual time sequences and an ML classifier for anomaly detection, CAM can produce visual explanations highlighting model attributions of its decisions.
-In our “Behavior Summarization” phase, we compute the similarity metrics of the original TS data and the model attributions, respectively, to conduct attribution-aware clustering from the latent space, as detailed in Sec.
-II-C.
-To enhance the efficiency of identifying model’s biased attribution issues, we further summarize each cluster with an integrated visualization, enabling experts to verify the attribution correctness without inspecting individual sequences.
-Our following “Annotation and Propagation” phase involves a visual interface for user interactions, where users overview clusters’ summarized model attributions, validate the exhibition of model issues, and annotate verified issues with a few clicks.
-As long as at least one cluster with “spurious attention” has been annotated, our backend method propagates the user’s assessment across other clusters as the estimated “Spuriousness scores”, enabling a facilitated model validation.
-Detailed explanations of this process can be found in Sec.
-II-D.
-After identifying the incorrect model’s attention, our model regularization strategy at the “Model Enhancement” stage systematically addresses these inaccuracies, refining the model’s focus to improve its overall performance and reliability, as described in Sec.
-II-E.
-Figure 3: HILAD applied to the anomaly detection of a univariate time series classifier trained on the MSL dataset [26].
-A System Menu, enabling the selection of dataset, class, and visualization options (Scatter plot view or Detailed plot view).
-B Cluster Information window, showing the data clusters and cluster metrics.
-C Representation Space window, showing a visual overview and the relative positions of all data clusters.
-D Instance Information window, showing individual time series sequences with model attributions belonging to a selected data cluster.
-II-B System Interface
-The visual interface of HILAD comprises four main components. Fig. 3 depicts our interface when validating an anomaly classifier on the Mars Science Laboratory (MSL) dataset.
+  3. A human-AI-collaborative solution for generating actionable insights for model validation and enhancement, which can inspire future research on human-assisted trustworthy AI.
+2 Related Works
+2.1 Time Series Anomaly Detection Models
+ML techniques for time series anomaly detection can be broadly categorized into unsupervised and supervised approaches.
+When ground-truth anomaly labels are unknown, unsupervised learning techniques are introduced to approximate future data points Kieu et al.
+(2019); Malhotra et al.
+(2015); Taylor and Letham (2018); Tuli et al.
+(2022), where the deviations between the predicted and true values serve as anomaly indicators.
+On the other hand, if labeled data are available, supervised models can be trained as classifiers to directly predict anomaly labels Del Campo et al.
+(2021); Gallicchio et al.
+(2017); Zhao et al.
+(2017).
+Such models can achieve higher accuracy and be utilized in critical areas like fraud detection or medical diagnostics Nabrawi and Alanazi (2023); Meena and Roy (2022).
+However, supervised models are generally more susceptible to bias issues, causing significant risks in practice Uddin et al.
+(2019).
+In our work, we focus on enhancing supervised models that directly predict anomaly labels, where various architectures such as transformers and CNN Dosovitskiy et al.
+(2020); Liu et al.
+(2021a); Boniol et al.
+(2023); Liu et al.
+(2021b); Yang et al.
+(2021) are fully supported, and we aim to detect, understand, and mitigate hidden biases in such models with a human-in-loop approach.
+2.2 Reliability Considerations of Models
+Biased models can cause significant risks in real-world applications Jung et al.
+(2021); Wang et al.
+(2024).
+For instance, if a model identifies normal time series as anomaly indicators, it underscores a fundamental misunderstanding of the underlying anomalous features, leading to potential overlooking of real anomalies Goswami et al.
+(2022); Deng and Kong (2021).
+While visual explanations such as CAM can reveal such discrepancies, their occurrence raises critical questions Xuan et al.
+(2024b) — Are these discrepancies a common feature across the entire data corpus, or are they isolated incidents?
+Moreover, once such biases are identified, what are the most effective methods for addressing and mitigating them?
+The discussion in Sec.
+1 points out the limitations of current XAI techniques in fully tackling these challenges.
+This scenario necessitates further research and development of methods that not only highlight but also rectify these biases.
+Enhancing the reliability of anomaly detection models through improved understanding and correction of biases is essential to foster trust and ensure the efficacy of these systems in real-world applications.
+2.3 Human-AI Collaboration
+The growing field of human-AI collaboration has demonstrated the importance of integrating human expertise into machine learning workflows, particularly in high-stakes applications such as medical imaging, finance, and industrial monitoring Amershi et al.
+(2019); Schaekermann et al.
+(2020); Xuan et al.
+(2022); Zhang et al.
+(2023); Yildirim et al.
+(2024); Xuan et al.
+(2024b); Deng et al.
+(2025).
+Prior research in interactive AI systems has explored how visual analytics (VA) platforms can support experts in understanding, validating, and improving models, with much of the focus on vision and natural language processing tasks Yang et al.
+(2022, 2023); Xuan et al.
+(2025a, b); Li et al.
+(2022a); He et al.
+(2021); Gou et al.
+(2020); Liu et al.
+(2012), as these tasks are generally more interpretable to humans compared to time series data.
+For example, VA systems designed in Xuan et al.
+(2025a); Li et al.
+(2022a) allow users to iteratively refine and probe deep learning models.
+Similar approaches in object detection and segmentation Gou et al.
+(2020); He et al.
+(2021); Yan et al.
+(2025) and medical AI Xuan et al.
+(2024b); Malik et al.
+(2025) have enabled human-guided model adjustments.
+These works underscore the potential of bidirectional human-AI collaboration, where users not only interpret model outputs but actively refine them through feedback loops.
+In time series anomaly detection, existing approaches Geiger et al.
+(2020); Alnegheimish et al.
+(2022); Liu et al.
+(2022); Feng et al.
+(2023) facilitate interactive anomaly investigation by displaying feature attributions and model decisions.
+While these methods improve interpretability, they often rely on instance-level explanations, limiting their scalability to larger datasets Ruta et al.
+(2019).
+As a result, users must perform labor-intensive manual inspections, which are impractical in complex industrial or financial environments where anomalies are highly dynamic and demand rapid adaptation.
+To address these challenges, we introduce HILAD to bridge the gap between visual interpretation and direct model refinement.
+HILAD supports a bidirectional interaction paradigm, where users can detect, interpret, and correct model misbehaviors at scale.
+Inspired by prior works that leveraged 2D projections for visualizing time series data Guo et al.
+(2021); Fujiwara et al.
+(2020), we incorporate a projection step to facilitate human interaction.
+However, unlike directly projecting raw features, we construct an attribution-aware latent space using an aggregated distance matrix over data and model attributions.
+HILAD leverages this space as an intermediate structure, supporting scalable annotation propagation and downstream model enhancement, which advances human-AI collaboration beyond prior visual-inspection-based approaches.
+By summarizing dominant failure patterns and allowing human adjustments, our system advances the integration of human expertise into AI-driven anomaly detection, enhancing model reliability and interpretability in time series applications.
+3 HILAD Framework
+Figure 2: The HILAD framework involves three phases: “Behavior Summarization”, “Annotation and Propagation”, and“Model Enhancement”.
+At the “Behavior Summarization” phase, we conduct attribution-aware clustering, where each cluster includes TS instances exhibiting similar data features and model attributions.
+We also visualize the overall pattern of each cluster to streamline user interpretation.
+At the “Annotation and Propagation” phase, users interact with our designed VA interface to identify and annotate model bias issues.
+These annotations are propagated to unannotated clusters to generate “Spuriousness scores” for user verification.
+Lastly, we mitigate the detected issues at the “Model Enhancement” phase.
+HILAD is a human-in-the-loop framework designed to improve the reliability of anomaly detection models on time series (TS) data.
+It takes as input a trained anomaly detection model along with TS instances with ground-truth anomaly labels.
+The system then enables domain experts to interactively inspect, validate, and correct model behaviors through three coordinated phases: “Behavior Summarization”, “Annotation and Propagation”, and “Model Enhancement”, as shown in Fig.
+2.
+The core idea of HILAD is to assist users in efficiently detecting and addressing model misbehavior caused by spurious attributions, where the model relies on misleading or irrelevant signals to make decisions.
+To support this, in the first “Behavior Summarization” phase, HILAD automatically generates and visualizes clusters with similar data features and model attributions.
+As illustrated in Fig.
+2, each cluster represents a group of TS sequences that share both contextual similarity in terms of data signals and model attribution similarity, and such patterns are visualized in “cluster summarization”, enabling experts to verify the attribution correctness without inspecting individual sequences.
+Then, in the following “Annotation and Propagation” phase, we enable users to overview clusters’ summarized model attributions, validate the exhibition of model issues, and annotate verified issues with a few clicks, where we provide label propagation to facilitate the validation process.
+Finally, the “Model Enhancement” stage systematically addresses these inaccuracies, refining the model’s focus to improve its overall performance and reliability.
+3.1 System Interface
+The visual interface of HILAD comprises four main components. Fig. 3 depicts our interface when validating an anomaly classifier on the Mars Science Laboratory (MSL) dataset Hundman et al. (2018).
 The System Menu (Fig.
@@
 The available metrics, including Accuracy, Confidence, and Spuriousness — the latter derived via the Spuriousness Propagation method (Sec.
-II-E) contingent upon user annotation — are presented to enhance the rapidity of analysis.
+3.4) contingent upon user annotation — are presented to enhance the rapidity of analysis.
 Once click/annotate one cluster in Cluster Information window, its corresponding position in 2D space will be highlighted in the Representation Space window (Fig.
@@
 We also provide information on the correctness of the current model in terms of classification and attention.
+Figure 3: HILAD applied to the anomaly detection of a univariate time series classifier trained on the MSL dataset Hundman et al.
+(2018).
+A System Menu, enabling the selection of dataset, class, and visualization options (Scatter plot view or Detailed plot view).
+B Cluster Information window, showing the data clusters and cluster metrics.
+C Representation Space window, showing a visual overview and the relative positions of all data clusters.
+D Instance Information window, showing individual time series sequences with model attributions belonging to a selected data cluster.
 Note that when interacting with the Representation Space window, the other two windows will be updated with coordinate information simultaneously to support user interpretation and issue verification.
@@
 Subsequently, our backend algorithm will fine-tune the model in accordance with the user’s annotation, followed by an update to the model’s performance metrics shown in System Menu.
-II-C Behavior Summarization
-At the “Behavior Summarization” phase, we first apply CAM to obtain model attributions correlated to its decision for each individual sequence.
-Then, we measure the similarities between: (1) the time-series data, according to the Dynamic Time Warping (DTW) Metric, and (2) the model attribution masks, according to the cosine similarity, respectively.
+3.2 Behavior Summarization
+At the “Behavior Summarization” phase, for each individual instance, we first apply CAM to obtain model attributions correlated to its decision, providing a one-dimensional vector (or mask) with the same length as the time sequence.
+Each value in this vector falls within the range (0, 1) and reflects the importance of the corresponding timestamp to the model’s prediction.
+Then, we measure and aggregate similarities between: (1) TS data, according to the Dynamic Time Warping (DTW) Metric, and (2) model attribution masks, according to the cosine similarity, respectively.
 With this information, we perform attribution-aware clustering to obtain clusters with similar data features and model attributions.
 To allow for effective user investigation, we visually summarize each cluster to help users identify common patterns of each cluster at a simple glance.
-We describe the details in the following.
-II-C1 Time-series Context Similarity Calculation
-In the realm of time series analysis, assessing the similarity between sequences is a pivotal task.
-A prevalent technique to measure this similarity is the DTW algorithm, which strives to minimize the distance between two time-sequenced datasets by dynamically aligning and “warping” them in the time dimension.
+Note that CAM is fully modular and can be replaced with a variety of attribution methods, including Grad-CAM Selvaraju et al.
+(2020), ScoreCAM Wang et al.
+(2020), and model-agnostic techniques like RISE Petsiuk et al.
+(2018).
+Our design does not rely on CAM-specific internals, allowing users to plug in alternative attribution mechanisms depending on the accessibility or specific requirements.
+We use CAM as our default XAI method, as CAM and CAM-derived techniques remain among the most widely used in time series and other domains due to their simplicity, interpretability, and computational efficiency.
+And they have been extensively validated across a broad range of models and datasets.
+3.2.1 Time series Context Similarity Calculation
+To measure the similarity between time sequences, we use the prevalent technique, the DTW algorithm, which strives to minimize the distance between two time-sequenced datasets by dynamically aligning and “warping" them in the time dimension.
 A smaller DTW distance indicates a higher similarity.
@@
 The initialization conditions are $D({T_{q}}_{n},{T_{p}}_{0})=D({T_{q}}_{0},{T_{p}}_{m})=\infty$ and $D({T_{q}}_{0},{T_{p}}_{0})=0$.
-II-C2 Attribution Similarity Calculation
-Given a fully convolutional network (FCN), $M$, trained for time series binary classification task (i.e., normal or abnormal), an input time series sequence, $T_{Q}$, and the predicted class of the $M$ for input $T_{Q}$, $c(T_{Q})$, the attention score of $T_{Q}$ is calculated as
-$a(T_{Q})=CAM(M,T_{Q},c(T_{Q})),$ (2)
-where $a(T_{Q})$ is a vector with the same dimension as $T_{Q}$, and each position in $a(T_{Q})$ corresponds to the degree of dependence of model $M$ on the corresponding dimension and time point when it makes the prediction $c(T_{Q})$.
-Thus $a(T_{Q})$ represents the model attribution.
-Please refer [21] to the specific calculation process of CAM.
-For input sequences $T_{Q}$ and $T_{P}$, the attribution similarity, $d_{cos}(a(T_{Q}),a(T_{P}))$, is computed using the cosine distance as:
+3.2.2 Attribution Similarity Calculation
+Given a model $\mathcal{M}$, trained for TS binary classification task (i.e., normal or abnormal), an input time sequence, $T_{Q}$, and the predicted class of the $\mathcal{M}$ for input $T_{Q}$, $c(T_{Q})$, the attention score of $T_{Q}$ is calculated as
+$a(T_{Q})=CAM(\mathcal{M},T_{Q},c(T_{Q})),$ (2)
+where $a(T_{Q})$ is the model attribution mask, which is a one-dimensional vector with the same length as $T_{Q}$.
+For input sequences $T_{Q}$ and $T_{P}$, the attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$ is computed by:
 $d_{cos}(a(T_{Q}),a(T_{P}))=1-\frac{a(T_{Q})\cdot a(T_{P})}{\|a(T_{Q})\|_{2}\times\|a(T_{P})\|_{2}},$ (3)
 where $\|a(T_{Q})\|_{2}$ and $\|a(T_{P})\|_{2}$ denote the L2 norm of vectors $a(T_{Q})$ and $T_{a}(P)$, respectively.
-Consistent with $d_{DTW}$, a smaller cosine distance indicates a higher similarity.
-II-C3 Attribution-Aware Clustering and Summarization
-To aggregate data with similar context and also similar model attribution together, we construct the following distance $m_{QP}$ to measure the similarity between two instances $T_{Q}$ and $T_{P}$:
+For this cosine distance, similar to $d_{DTW}$, a smaller value indicates a higher similarity.
+3.2.3 Attribution-Aware Clustering and Summarization
+To aggregate data with both similar context and similar model attribution, we construct the following aggregated distance $m_{QP}$ to measure the similarity between two instances $T_{Q}$ and $T_{P}$:
 $m_{QP}=\alpha d_{DTW}(T_{Q},T_{P})+(1-\alpha)d_{cos}(a(T_{Q}),a(T_{P})),$ (4)
 where $\alpha$ is the weight to balance the data context similarity $d_{DTW}(T_{Q},T_{P})$ and the model attribution similarity $d_{cos}(a(T_{Q}),a(T_{P}))$.
-In this work, we set $\alpha=0.5$ in all case studies.
-Thus, for all of the $N$ time series instances, we can get the following distance matrix:
-$M_{dist}=\begin{bmatrix}m_{1,1}&m_{1,2}&\cdots&m_{1,N}\\ m_{2,1}&m_{2,2}&\cdots&m_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ m_{N,1}&m_{N,2}&\cdots&m_{N,N}\end{bmatrix}.$ (5)
-Finally, we use UMAP [27] to plot $M_{dist}$ and apply K-means [28] to generate $K$ clusters:
+Both distances are normalized to the range $[0,1]$ to ensure scale compatibility and prevent either term from dominating the aggregated distance.
+In this work, we set $\alpha=0.5$.
+Thus, for all of the $N$ TS instances, we can get a distance matrix $M_{dist}\in\mathbb{R}^{N\times N}$.
+Finally, we use UMAP McInnes et al. (2018) to project $M_{dist}$ and apply K-means Krishna and Murty (1999) to generate $K$ clusters:
 $C=\text{K-means}(UMAP(M_{dist}),K),$ (5)
-where $C=\{c_{1},c_{2},\dots,c_{K}\}$ is the set of clusters, UMAP$(M_{dist})$ is the 2D representation of the data and $K$ is the desired number of clusters.
+where $C=\{c_{1},c_{2},\dots,c_{K}\}$ is the set of clusters, UMAP$(M_{dist})$ is the 2D representation of the data, and $K$ is the desired number of clusters determined by the elbow method.
 As a follow-up step, we visually summarize the resulting clusters as an integrated visualization for each cluster, which presents the common patterns of time sequences and model attributions in the current cluster, empowering users to grasp model behavior patterns more efficiently.
-II-D Annotation and Propagation
-Building upon the aforementioned method, at the “Annotation and Propagation” phase, we To effectively evaluate such spurious attention, we introduce a metric termed “Spuriousness score”, $S[c_{i}]$, ranging from 0 to 1, signifying the probability range from the least to the most likelihood of spurious attention occurrence.
-Furthermore, we deploy the Label Propagation [29], which automatically estimates probabilities of the unannotated clusters’ spuriousness based on users’ annotations and neighboring clusters $N(c_{i})$ in the representation space.
+3.3 Annotation and Propagation
+Building upon the aforementioned method, at the “Annotation and Propagation” phase, we define a metric termed “Spuriousness score", $S[c_{i}]$, to evaluate the presence of spuriousness.
+This score ranges from 0 to 1, representing the spurious probability from the least to most likely.
+Furthermore, the Label Propagation algorithm Zhou et al.
+(2003) is deployed to automatically estimate probabilities of the unannotated clusters’ spuriousness based on users’ annotations and neighboring clusters $N(c_{i})$.
 The representation space is generated by UMAP as described in Sec.
-II-C, where we use the adjacency matrix, $A(i,j)$, to denote the distance between clusters’ feature representations of $c_{i}$ and $c_{j}$.
-Notably, users only need to provide an annotation for one cluster (Algo.
-1, Line 6), designating it as either “correct attention” ($S[c_{i}]=0$) or “spurious attention” ($S[c_{i}]=1$).
+3.2, where we use the adjacency matrix, $A(i,j)$, to denote the distance between clusters’ feature representations of $c_{i}$ and $c_{j}$.
+Notably, users only need to provide binary annotation for a cluster, designating it as either “correct attention" ($S[c_{i}]=0$) or “spurious attention" ($S[c_{i}]=1$).
 Our algorithm (Algo.
-1, Line 8-15) then propagates this information to compute Spuriousness scores for other clusters.
+1, Line 6-14) then propagates this information to compute Spuriousness scores for other clusters.
 Subsequently, these scores are showcased in the Cluster Information window (Fig.
 3  B).
-The “Spuriousness score” yields two noteworthy advantages.
-Primarily, they facilitate a more streamlined exploration process of clusters, with potentially spurious correlations being conspicuously highlighted.
-This is a pivotal step in aiding users to identify and assess problematic attention.
-Secondly, following users’ validation, these scores are leveraged to determine the problematic clusters crucial for issue mitigation, as we elucidate in the following section.
 Algorithm 1 Label Propagation for Spuriousness Score
@@
 3:$A$: Adjacency matrix of clusters indicating similarity.
-4:
-5:$\mathbf{S}$ : Vector of spuriousness scores for each cluster.
-6:Initialize $\mathbf{S}$ with -1 for all clusters.
-7:for each cluster $c_{i}$ in $C$ with a user label do
-8:     $S[c_{i}]$ = label value of $c_{i}$
-9:while unlabeled clusters exist in $C$ do
-10:     for each unlabeled cluster $c_{i}$ in $C$ do
-11:         $\text{WeightedSum}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})\times S[c_{j}]$
-12:         $\text{TotalWeight}(c_{i})=\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})$
-13:         $S[c_{i}]=\frac{\text{WeightedSum}(c_{i})}{\text{TotalWeight}(c_{i})}$
-14:return $\mathbf{S}$
-II-E Model Enhancement
-An effective strategy for addressing detected spurious correlations in machine learning (ML) models involves model retraining or finetuning, which reduces the model’s resilience over potential bias while keeping the architecture intact.
-Retraining the model on a noise-corrupted dataset has demonstrated its efficacy in diminishing the model’s reliance on spurious features [30].
-To counteract the influence of spurious features, we harness the masking strategy to shift and correct the model’s attention.
-The idea is conceptually simple: first, mask the parts of the data that receive spurious attention, then encourage the model to infer how to make decisions with the rest of the data.
-This idea has been used as a self-supervised pre-training method and spuriousness feature elimination method in Computer Vision (CV) and Natural Language Processing (NLP) fields.
-In our work, we adjust the Core Risk Minimization (CoRM) methodology [30] in the CV domain into time series data to introduce controlled masks into the annotated spurious attention time points by injecting random Gaussian noise into the data.
-This is also a fundamental data augmentation strategy called jittering for time series data that can help the model become more robust to variations and be used in masked autoencoder for time series forecasting task [31].
+4:$\mathbf{Y}$: Initial labels, where $Y[c_{i}]\in\{0,1\}$ for labeled clusters and $-1$ otherwise.
+5:
+6:$\mathbf{S}$: Vector of propagated spuriousness scores for each cluster.
+7:Initialize $\mathbf{S}\leftarrow\mathbf{Y}$
+8:for iteration = 1 to max_iter do
+9:     $\mathbf{S}^{\text{new}}\leftarrow\mathbf{S}$
+10:     for each cluster $c_{i}\in C$ do
+11:         if  $Y[c_{i}]\neq-1$  then
+12:              continue
+13:         $\text{WeightedSum}\leftarrow\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})\times S[c_{j}]$
+14:         $\text{TotalWeight}\leftarrow\sum_{c_{j}\in N(c_{i})}A(c_{i},c_{j})$
+15:         $S^{\text{new}}[c_{i}]\leftarrow\frac{\text{WeightedSum}}{\text{TotalWeight}}$
+16:     $\mathbf{S}\leftarrow\mathbf{S}^{\text{new}}$
+17:return $\mathbf{S}$
+The “Spuriousness score" yields two noteworthy advantages.
+Primarily, they facilitate a more streamlined exploration process of clusters, with potentially spurious correlations being conspicuously highlighted.
+This is a pivotal step in aiding users to identify and assess problematic attention.
+Secondly, following users’ validation, these scores are leveraged to determine the problematic clusters crucial for issue mitigation, as we elucidate in the following section.
+3.4 Model Enhancement
+An effective strategy for addressing detected spurious correlations involves model retraining or finetuning, which reduces the model’s resilience over potential bias while keeping the architecture intact.
+Specifically, the Core Risk Minimization (CoRM) methodology Singla et al.
+(2022) that retrains the model on a noise-corrupted dataset has demonstrated its efficacy in spuriousness mitigation for vision models.
+In our work, we adapt CoRM to TS data, which harnesses the masking strategy by injecting random Gaussian noise into the spurious time points, aiming to shift and correct the model’s wrong attention by encouraging it to infer decisions with the rest of the data.
+The idea aligns with a data augmentation strategy called jittering for time series data, which can help the model become more robust to variations and be used in masked autoencoders for time series forecasting Tang and Zhang (2022).
 The integration of CoRM into HILAD entails several steps.
-First, after users annotate the data, we extract a set of instances, $T_{S}$, that display heightened probabilities of spuriousness, indicating undesirable attention.
-We also extract a set of instances, $T_{C}$, that display heightened probabilities of correctness, indicating correct attention.
-For the instances in $T_{S}$, pixel attributions (specifically, CAM masks) accentuate the spurious regions.
-We employ these masks to introduce random Gaussian noise to the identified spurious regions.
-Mathematically, for an individual time sequence $T_{i}\in T_{S}$, this operation can be succinctly represented as
+We first extract a set of instances with high spuriousness scores, $T_{S}$, indicating undesirable attention.
+We also extract a set of instances with the lowest spuriousness, $T_{C}$, indicating correct attention.
+The attribution masks of $T_{S}$ highlight spurious regions, which are used to occupy such regions with random Gaussian noise accordingly.
+For an individual time sequence $T_{i}\in T_{S}$, this operation can be mathematically represented as
 $T^{\prime}_{i}=T_{i}+\mathbf{m}\odot z,$ (6)
-wherein $T_{i}$ is the input time series data, $\mathbf{m}$ denotes the CAM mask, $z$ represents the generated Gaussian noise matrix, and $\odot$ denotes the Hadamard product.
-Each variable is of the same dimensions as the input time series data.
+where $\mathbf{m}$ denotes the CAM mask, $z$ represents the generated Gaussian noise matrix, and $\odot$ denotes the Hadamard product.
 Similarly, for $T_{j}\in T_{C}$, we add random Gaussian noise to the non-highlighted areas to enhance the model’s attention to abnormal patterns:
@@
 After substituting the original data with these spuriousness-masked and correctness-enhanced data, we proceed to retrain the model.
-Subsequently, an evaluation is conducted to determine the extent to which spurious correlations have been alleviated.
+Subsequently, an evaluation is conducted to measure the mitigation of spurious correlations.
 In Sec.
-III-D, we elaborate on the evaluation metrics utilized to quantify the effectiveness of our method in mitigating spurious correlations.
-III Experimental Design
+4.6, we elaborate on the evaluation metrics utilized to quantify the effectiveness of our method in mitigating spurious correlations.
+4 Experimental Design
 In this section, we provide an experimental design to benchmark and evaluate the capabilities of HILAD.
 The primary objective of this experiment is to investigate how HILAD empowers humans to detect, understand, and correct potential issues in an anomaly classifier for time series data.
-III-A Task
+4.1 Datasets
+Mars Science Laboratory (MSL) Dataset Hundman et al.
+(2018) corresponds to the sensor and actuator data for the Mars rover.
+The dataset provides the ground truth anomaly masks that marks the time point when the abnormal behavior occurs.
+Since this dataset is known to have many trivial sequences Wu and Keogh (2021), to make a best-faith effort to eliminate mislabeled time points, we build a new dataset based on the two nontrivial ones (A2 and A4).
+Specifically, we randomly divide the original time series into 1000 equal-length sequences, all of 800 lengths, where the ratio of normal to abnormal data is 3:2.
+Simulation Testbed for Exploration Vehicle ECLSS (STEVE) Deng et al.
+(2023) is a simplified single-bed $CO_{2}$ removal system of the Carbon Dioxide Removal Assembly (CDRA) onboard the International Space Station (ISS), at the University of Colorado Boulder.
+A corresponding Simulink model is provided to simulate multiple failure modes similar to the STEVE testbed.
+In our study, we conducted a simulation consisting of four cycles, each comprising 80 minutes of $CO_{2}$ adsorption followed by 80 minutes of desorption.
+During the third cycle, a simulated leak failure is introduced at a random time point in Valve 1, resulting in observable anomalies in three sensors: Bed Temperature, $CO_{2}$ Concentration, and Flow Rate.
+These sensor sequences are treated collectively.
+We generated 1000 sets of sequences with a ratio of 4:1 between normal and abnormal data.
+Additionally, we marked time points at which abnormal behaviors occurred in the three sensors as ground truth masks.
+KPI anomaly detection for AIOps (KPI) dataset Li et al.
+(2022b) is a real-world, large-scale dataset for evaluating anomaly detection systems in industrial AIOps settings.
+It contains time series with noisy, high-variance patterns and diverse anomaly types, making it suitable for benchmarking real-world deployment performance.
+For our experiments, we construct equal-length sequences of 800 time points from four KPI categories with the highest anomaly ratios, resulting in 1,200 time series instances.
+Additional dataset details are provided in the Appendix.
+4.2 Baselines
+Baseline models.
+We involve two baseline models in our experiments.
+One is a fully convolutional network (FCN) Wang et al.
+(2017), which has demonstrated effectiveness on time series classification.
+The other is TST Zerveas et al.
+(2021), a transformer-based model for multivariate time series anomaly detection.
+These two models have distinct architectures for evaluating our method with various baselines.
+Note that HILAD is fully applicable to other time series anomaly classification models with different architectures Boniol et al.
+(2023); Liu et al.
+(2021b); Yang et al.
+(2021).
+In addition to baseline models, we compare HILAD with the following spuriousness mitigation strategies using data augmentation:
+•  Random aug.   Randomly select $\Gamma$ instances from the training set for model fine-tuning.
+•  Random mask.   Select $\Gamma$ instances proposed by our system, but mask random time points.
+•  Cluster mask.
+The $\Gamma$ instances proposed by our system only concede the participant’s operation in the Cluster Information window, which means the instance-level operations are discarded.
+This method uses the same masking strategy as ours.
+4.3 Participants
+This experiment was approved by the Institutional Review Board.
+We involve 15 participants in total, of which 13 were male and 2 were female.
+All participants are graduate students between the ages of 23 to 30 years old.
+All participants passed the test to ensure they could accurately identify and annotate the abnormal pattern.
+4.4 User Tasks
 In the experiment, participants were asked to interact with the interface of HILAD to evaluate a classification model trained on a specific dataset, containing both normal and abnormal time sequences.
-Specifically, we encouraged them to use as many visual components as possible, follow the system guidance to verify the highlighted issues, interpret them with the provided information, and annotate the confirmed errors in a restricted time.
-In each experiment, participants were only allowed to perform a single model retraining to better control experiment time and quantify the evaluation results.
+Specifically, we encouraged them to use as many visual components as possible, follow the system guidance to verify the highlighted issues, interpret them with the provided information, and annotate the confirmed errors within a restricted time.
+In each experiment, participants were only allowed to perform model retraining once to better quantify the results.
 With the awareness of the HILAD workflow, participants explored the general model behavior concerning different clusters of data through the Cluster Information window, where they also investigated further details from the Instance Information window that updated simultaneously with their interactions.
-They were asked to mark and annotate detection issues, such as a model using normal time points to determine an abnormal time series, which could be achieved by simply clicking the “Spurious” button in the Cluster Information window.
+They were asked to mark and annotate detection issues, such as a model using normal time points to determine an abnormal time series, which could be achieved by simply clicking the “Spurious" button in the Cluster Information window.
 At this point, all instances corresponding to this cluster were marked as spurious.
@@
 As long as one cluster is annotated as spurious or correct, our algorithm will approximate the attention correctness of other clusters according to their similarities in the representation space, resulting in a spuriousness approximation score for every cluster.
-Participants can click the “spuriousness” button provided in our interface to rank clusters according to this score, allowing for adjustments and confirmation of the approximated model issues.
+Participants can click the “spuriousness" button provided in our interface to rank clusters according to this score, allowing for adjustments and confirmation of the approximated model issues.
 When the participants confirmed the first round of issue inspection could be concluded, they would click the retrain button, and our system performed the model retraining based on the participants’ feedback.
-III-B Procedure
-In the following, we introduce the experiment’s procedure. The overall practical accomplishment of the experiment took between 30 and 45 minutes.
-III-B1 Introduction and Preparations
+4.5 Procedure
+In the following, we introduce the procedure of our user study.
+4.5.1 Introduction and Preparations
 Before the experiment begins, participants receive detailed instructions explaining the experiment’s purpose and overall structure before it begins.
-Next, they are introduced to the key concepts involved in time-series classification tasks, including metrics such as confidence and accuracy, possible abnormal patterns, and spurious model attention.
+Next, they are introduced to the key concepts involved in time series classification tasks, including metrics such as confidence and accuracy, possible abnormal patterns, and spurious model attention.
 Subsequently, an in-depth orientation to HILAD is provided, highlighting its various components and functionalities.
@@
 The primary goal communicated to the participants is to employ HILAD for monitoring effectively and analyzing the model’s behavior, focusing on its attention mechanisms.
-III-B2 Familiarization Procedure
+4.5.2 Familiarization Procedure
 A simplified dataset was introduced to acquaint participants with the functions of the HILAD interface.
@@
 Once participants demonstrated proficiency in accurately identifying and annotating abnormal patterns, they conducted the experimental run.
-III-B3 Experimental Run
-In the experimental run, participants were introduced to the HILAD with two space domain-related datasets that mimic its real-world use for anomaly detection.
-Their task was to critically analyze the behavior of the classification model across the datasets.
-The main variable in this phase was the dataset itself, and the sequence of dataset presentation varied systematically among participants.
-This approach aimed to mitigate any potential bias from learning curve effects.
+4.5.3 Experimental Run
+In the experimental run, participants’ tasks were to critically analyze the behavior of each classification model across three datasets.
 We executed three trials for each dataset to further effectively compare the distinctions between the automated algorithm and the human-AI collaboration method.
-These trials varied in the scope of information provided to the participants: the first trial provided access to only the System Menu window, the second trial included System Menu, Cluster Information, and Representation Space windows, and the third trial offered a comprehensive view with the complete all of the four windows.
+These trials varied in the scope of information provided to the participants: the first trial provided access to only the System Menu window, the second trial included System Menu, Cluster Information, and Representation Space windows, and the third trial offered a comprehensive view with the complete all four windows.
 The first trial in our study exemplifies the application of the automated algorithm, while the latter two trials represent the human-AI collaboration methods.
@@
 This approach was essential to assess the impact of varying levels of information availability on the performance and decision-making processes in the context of human-AI interaction.
-III-B4 Postprocessing
+4.5.4 Postprocessing
 Once all experimental runs are concluded, participants can reflect on the experiment and share their insights and feedback regarding HILAD.
@@
 This data provides valuable insights into the user experience and the system’s effectiveness in facilitating anomaly detection, evaluation, and correction.
-III-C Participants
-This experiment was approved by the Institutional Review Board at the University of California, Davis.
-We involve 15 participants in total, of which 13 were male and 2 were female.
-All participants are graduate students between the ages of 23 to 30 years old.
-All participants passed the test to ensure they could accurately identify and annotate the abnormal pattern.
-All completed the entire experiment on two provided datasets.
-III-D Performance Evaluation Metrics
+4.6 Performance Evaluation Metrics
 We measure the effectiveness of our proposed framework in terms of classification and attention performance.
-From the perspective of classification performance, we use accuracy ($Acc$), precision ($P$), recall ($R$), and $F_{1}$ score to evaluate the performance of our method comprehensively.
-The details of the metrics are listed as follows:
-$Acc=\frac{TP+TN}{TP+TN+FP+FN},P=\frac{TP}{TP+FP},$
-$R=\frac{TP}{TP+FN},F_{1}=\frac{2\cdot P\cdot R}{P+R},$
-where TP, TN, FP, and FN are four classification results, denoting true positive, true negative, false positive, and false negative, respectively.
-$Acc$ is a measure of the overall correctness of a classification model.
-$P$ represents the ability of the method to distinguish anomalies, and a higher $P$ accurately indicates fewer false anomalies.
-$R$ is to determine the ability of the method to detect all the anomalies.
-Since The $P$ and $R$ metrics often appear contradictory, the $F_{1}$ score is a metric for comprehensively considering these two indicators and providing a balanced overall performance of the method.
-When the data set has imbalanced classes, F1 Score might be a better metric than accuracy.
-In the perspective of attention accuracy performance, we adjust the relevance accuracy ($RA$) proposed in [23] to quantify how accurate a model is at locating ground-truth time points at which the abnormal behavior occurs.
+From the perspective of classification performance, we use the standard Accuracy, Precision, Recall, and $F_{1}$ Score to evaluate the performance of our method comprehensively.
+In the perspective of attention accuracy performance, we adjust the relevance accuracy ($RA$) proposed in Wickstrøm et al.
+(2020) to quantify how accurate a model is at locating ground-truth time points at which the abnormal behavior occurs.
 The core idea of relevance accuracy is to compare the $k$ most attention time points for a prediction with the ground-truth abnormal time points.
@@
 The model needs to highlight some time points as being more important than others.
-To assess the participants’ subjective perception of human–machine cooperation, a questionnaire with a five-point Likert Scale [24] with the following items was used.
-Figure 4: Tukey HSD post-hoc test results for MSL dataset.
-The left plot illustrates classification performance, while the right one depicts attention accuracy.
-Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
-III-E Datasets
-Mars Science Laboratory (MSL) dataset is a dataset that corresponds to the sensor and actuator data for the Mars rover [26].
-At the same time, the data set marks the time point when the abnormal behavior occurs for us to calculate relevance accuracy.
-However, this dataset is known to have many trivial sequences [32]; hence, to make a best-faith effort to filter out data with mislabeled ground truth time points, we build the new dataset based on the two nontrivial ones (A2 and A4), where we randomly divide the original time series into 1000 equal-length sequences, all of 800 lengths, where the ratio of labeled normal to abnormal data is 3 to 2.
-The Simulation Testbed for Exploration Vehicle ECLSS (STEVE) [33] at the University of Colorado Boulder is a simplified single-bed $CO_{2}$ removal system of the Carbon Dioxide Removal Assembly (CDRA) onboard the International Space Station (ISS).
-A corresponding Simulink model is provided to simulate multiple failure modes similar to the STEVE testbed.
-In our study, we conducted a simulation consisting of four cycles, each comprising 80 minutes of $CO_{2}$ adsorption followed by 80 minutes of desorption.
-During the third cycle, a simulated leak failure is introduced at a random time point in Valve 1, resulting in observable anomalies in three sensors: Bed Temperature, $CO_{2}$ Concentration, and Flow Rate.
-These sensor sequences are treated collectively as a set of sequences.
-We generated 1000 sets of sequences with a ratio of 4 to 1 between labeled normal and abnormal data.
-Additionally, we marked the time points at which abnormal behaviors occurred in the three sensors to calculate relevance accuracy.
-III-F Baseline Methods
-Baseline model.
-In this paper, we use the fully convolutional network (FCN) proposed in [34], which has demonstrated good performance on time series classification benchmarks.
-In addition to the baseline model, we also compare our spurious attention elimination method with the following data augmentation strategies for the same number, $Q$, of instances:
-Random Aug. Randomly select $Q$ instances from the training set for model fine-tuning.
-Random Mask.
-Select the $Q$ instances proposed by our system, but add masks at random time points for each instance.
-To further verify the effectiveness of our method, we compared the following strategy as the ablation study of our method.
-Cluster Mask.
-The $Q$ instances proposed by our system only concede the participant’s operation in the Cluster Information window, which means the participants in the Instance Information window make no further adjustments.
-This method uses the same masking strategy as ours.
-III-G Statistical Analysis
-For Hypothesis 1, since all five methods involved in the comparison have been tested an equal number of times, we directly assess their performance using the metrics introduced in Section III-D.
+To assess the participants’ subjective perception of human–machine cooperation, a questionnaire with a five-point Likert Scale De Winter and Dodou (2010) with the following items was used.
+4.7 Statistical Analysis
+For Hypothesis 1, since all five methods involved in the comparison have been tested an equal number of times, we directly assess their performance using the metrics introduced in Section 4.6.
 Therefore, our hypothesis tests reduces to:
@@
 Conversely, by contrasting (2) with (3), we seek to discern the added benefits by enabling detailed individual data adjustments within the clusters.
-IV Results
-Figure 5: Tukey HSD post-hoc test results for STEVE dataset.
-The left plot illustrates classification performance, while the right one depicts attention accuracy.
+4.8 Annotation Efficiency Evaluation Settings
+We quantify and compare the human effort between HILAD and multiple representative baselines that also utilize human annotations.
+Specifically, our comparative study involves three active learning baselines: random sampling, uncertainty-based sampling, and diversity-based sampling.
+To align with prior practices, we view our basic model as the warm-up model that is trained with labeled data and might have spurious issues.
+Then, during the human interaction phase, we present the user with individual time sequences with CAM highlighting, and ask them to select from binary options, “correct” or “spurious”, indicating whether spurious correlation exists.
+View D in the HILAD interface is used as the annotation interface.
+We fix the annotation budget as 160, which is 20% of the total instances.
+Additional settings are as follows.
+  • Random-based sampling. Instances are randomly sampled from the dataset and listed in the interface for user annotation. This reflects real-world, unstructured manual debugging workflows.
+  • Uncertainty-based sampling.
+Individual instances are ranked in descendingly order by model uncertainty (e.g., based on softmax entropy).
+Users are given instances in order and assign spuriousness labels to each of them.
+  • Diversity-based sampling. We use the basic setting to cluster data features and sample instances nearest to each cluster centroid. Then present the instances to the user in a random order.
+Considering the practical challenges, we re-invite five of our participants in the original study.
+We asked them to perform annotations following four settings: ours, random, uncertainty, and diversity.
+Their orders are random to reduce familiarity-caused noise, and we have a 10-minute break in between.
+Note that for HILAD, the users are asked to determine how many instances to annotate, according to their interactions and observations through our interface.
+5 Results
+Figure 4: Tukey HSD post-hoc test results for MSL dataset using FCN.
+The left and right plots illustrate classification and attention accuracy, respectively.
 Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
-Figure 6: Participants perception of the system, according to nine Likert-Type Questions using a 5-point scale.
-In order to effectively measure the objective performance difference between purely automatic algorithm-driven and human-machine cooperative anomaly detection methods, we conducted an analysis using Analysis of Variance (ANOVA) tests [35].
+Note: y-axes are truncated to emphasize fine-grained differences.
+Figure 5: Tukey HSD post-hoc test results for STEVE dataset using FCN.
+The left and right plots illustrate classification and attention accuracy, respectively.
+Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
+Note: y-axes are truncated to emphasize fine-grained differences.
+5.1 Quantitative evaluation
+In order to effectively measure the objective performance difference between purely automatic algorithm-driven and human-machine cooperative anomaly detection methods, we conducted an analysis using Analysis of Variance (ANOVA) tests St et al.
+(1989).
 Objective performance is divided into two aspects, namely classification correctness and attention correctness.
@@
 Upon identifying significant differences with the ANOVA test, we proceeded with the Tukey HSD post-hoc test to further distinguish which specific group means, among the compared groups, are significantly different from each other, ensuring a thorough evaluation of pairwise contrasts.
-Owing to space limitations, this article presents only the results of the Tukey HSD post-hoc test for two datasets, as illustrated in Fig.
-4 and Fig.
-5, respectively.
+We report the Tukey HSD post-hoc test results on the MSL and STEVE datasets using FCN in Figs 4 and 5.
+For the more challenging real-world KPI dataset, we include both FCN and TST models, with results shown in Figs.
+6 and 7.
+Across all settings, HILAD consistently outperforms baseline and augmentation-based methods in both prediction accuracy and attribution alignment.
+These results underscore the importance of evaluating not only classification outcomes but also the underlying rationale, reinforcing the need for human-in-the-loop validation.
+5.2 Annotation efficiency evaluation
+The results are reported in Table 1, which includes the average annotation time, annotated data size, the spuriousness label coverage (how many instances have been assigned a label), and the F1 gain in percentage compared to the basic model.
+We can observe that HILAD consistently achieved the best trade-off: requiring the least annotation time (8.3 $\pm$ 2.1 min), while delivering the highest F1 improvement (18.485 $\pm$ 2.7%) and covering over 800 sequences thanks to our propagation design.
+In contrast, despite much larger annotation budgets being used, the active learning baselines still underperformed because they solely rely on manual annotations, leading to significantly less spuriousness label coverage.
+These results demonstrate that our framework supports minimal but essential human input.
+With fewer human annotations, our design of attribution-guided behavior summarization and annotation propagation leverages human efforts at a lower cost, which is not achievable through unstructured or conventional data sampling.
+Figure 6: Tukey HSD post-hoc test results for KPI dataset using FCN.
+The left and right plots illustrate classification and attention accuracy, respectively.
+Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
+Note: y-axes are truncated to emphasize fine-grained differences.
+Figure 7: Tukey HSD post-hoc test results for KPI dataset using TST.
+The left and right plots illustrate classification and attention accuracy, respectively.
+Methods not sharing any letter are significantly different by the Tukey-test at the 5% level of significance.
+Note: y-axes are truncated to emphasize fine-grained differences.
+Method Avg. annotation time (minutes) Avg. annotation number Coverage F1 Gain (%)
+random 14.2 $\pm$ 1.8 160 160 1.908 $\pm$ 1.2
+uncertainty 13.3 $\pm$ 1.4 160 160 2.610 $\pm$ 0.4
+diversity 14.7 $\pm$ 2.3 160 160 2.328 $\pm$ 1.9
+HILAD (ours) 8.3 $\pm$ 2.1 10.2 $\pm$ 3.4 800 18.485 $\pm$ 2.7
+Table 1: Comparison of annotation time, number, coverage, and resulting F1 gain across HILAD and baseline annotation strategies on MSL dataset.
+5.3 Objective evaluation
 The result of the five-point Likert scale for objective evaluation is shown in Fig.
-6.
+8.
 For a more comprehensive analysis, we also furnish the post-test results for the three methods, assessing their comparative effectiveness using a t-test.
-The stars intended to flag levels of significance as shown in Fig.
-6 are based on the adjusted p-values.
-V Discussion
-V-A Hypothesis 1
+The stars intended to flag levels of significance, which are based on the adjusted p-values.
+5.4 Ablation study
+To validate our design choice of $\alpha$ = 0.5 in the joint similarity computation for behavior clustering, we conducted a sensitivity analysis by varying $\alpha$ across a range of values.
+Refer to Formula 4 in Sec.
+3.2.2, $\alpha$ controls the balance between time sequence similarity (via DTW) and attribution similarity (via CAM), with higher $\alpha$ favoring time sequence similarity.
+We first vary $\alpha$ across $\{0.0,0.5,1.0\}$ and provide qualitative observations on the produced clusters in Fig.
+9.
+Specifically, in each setting, we randomly sample three clusters from the results and observe their included instances.
+The results show $\alpha=0.0$ overlooks the data consistency and $\alpha=1.0$ contrarily misses the model attribution consistency, while our choice $\alpha=0.5$ produces clusters with internally consistent time sequences and model attributes simultaneously.
+Figure 8: Participants perception of the system, according to Likert-Type Questions using a 5-point scale.
+We also quantitatively evaluate the impact of $\alpha$ on model performance, where we utilize FCN as the baseline on MSL data, measuring classification performance and attention correctness.
+In the experiment, we only allow cluster-level annotation and disable the adjustment of individual instances to ensure valid evaluation of $\alpha$.
+As shown in Table 2, $\alpha=0.5$ consistently produces the highest F1 and attention scores, supporting the use of equal weighting between sequence context and model behavior.
+Extreme values (e.g., $\alpha=0.0$ or $1.0$) led to performance drops, where $\alpha=1.0$ (only consider data similarity) results in the worst performance, as we lost neighbor consistency regarding model attributions.
+This experiment validates our parameter choice, confirming that collectively considering both temporal patterns and model attention is critical for meaningful behavior grouping.
+6 Discussion
+6.1 Hypothesis 1
 Firstly, regarding classification correctness, Fig.
@@
 This observation reinforces the notion that our method not only objectively enhances classification performance but also elevates attention accuracy, thereby bolstering overall system reliability.
+Figure 9: Qualitative observation on randomly sampled clusters when using different $\alpha$ values.
+We can observe that $\alpha=0$ leads to clusters with more coherent model attention only, while $\alpha=1$ produces clusters with consistent time sequences but dissimilar model behavior patterns.
+In contrast, our choice of $\alpha=0.5$ provides clusters with both consistent model attribution and similar time sequences.
+$\alpha$=0.00 $\alpha$=0.25 $\alpha$=0.50 $\alpha$=0.75 $\alpha$=1.0
+F1 88.21 $\pm$ 0.26 89.92 $\pm$ 0.27 90.62 $\pm$ 0.25 88.33 $\pm$ 0.60 87.1 $\pm$ 0.76
+ATT 86.43 $\pm$ 0.79 89.13 $\pm$ 0.78 89.90 $\pm$ 0.77 85.82 $\pm$ 0.85 85.66 $\pm$ 0.84
+Table 2: Quantitative evaluation on the effect of the $\alpha$ parameter that controls the balance between temporal and attribution similarity in clustering, showing $\alpha=0.5$ achieves the best trade-off in both classification (F1) and attention (ATT) performance.
 In summary, our method exhibits substantial deviations in both classification correctness and attention correctness when compared to the automatic algorithm across the two datasets.
 This substantiates our conjecture and underscores the empirical evidence supporting our approach.
-V-B Hypothesis 2
+6.2 Hypothesis 2
 In the subjective evaluation, both the cluster mask method and our method are significantly different from the automatic algorithm in intuition, transparency, and reliability.
-Because the automatic algorithm can only provide the result of classification, can not provide the criterion of classification can not interact.
+Because the automatic algorithm only provides classification results, it neither reveals the underlying decision criteria nor supports interaction.
 The cluster mask method is not significantly different from intuition and transparency, as both provide reasons for the model to make a prediction.
@@
 Further, on the objective and subjective level, the necessity of introducing D window is verified by comparing with Cluster Mask method.
-VI Conclusion
+7 Conclusion
 In conclusion, we introduce a human-AI collaboration framework, HILAD, to enhance the performance and reliability of time series anomaly detection models.